# Quantization Meets dLLMs: A Systematic Study of Post-training Quantization for Diffusion LLMs

### v1

@@ Abstract
Recent advances in diffusion large language models (dLLMs) have introduced a promising alternative to autoregressive (AR) LLMs for natural language generation tasks, leveraging full attention and denoising-based decoding strategies. However, the deployment of these models on edge devices remains challenging due to their massive parameter scale and high resource demands. While post-training quantization (PTQ) has emerged as a widely adopted technique for compressing AR LLMs, its applicability to dLLMs remains largely unexplored. In this work, we present the first systematic study on quantizing diffusion-based language models. We begin by identifying the presence of activation outliers, characterized by abnormally large activation values that dominate the dynamic range. These outliers pose a key challenge to low-bit quantization, as they make it difficult to preserve precision for the majority of values. More importantly, we implement state-of-the-art PTQ methods and conduct a comprehensive evaluation across multiple task types and model variants. Our analysis is structured along four key dimensions: bit-width, quantization method, task category, and model type. Through this multi-perspective evaluation, we offer practical insights into the quantization behavior of dLLMs under different configurations. We hope our findings provide a foundation for future research in efficient dLLM deployment. All codes and experimental setups will be released to support the community.
@@ 1 Introduction
Current efforts toward optimizing dLLM inference have primarily focused on designing specialized key-value (KV) cache mechanisms  (Wu et al., 2025; Ma et al., 2025; Liu et al., 2025b; Wang et al., 2025). However, quantization  (Li et al., 2024; Liu et al., 2025a; Wei et al., 2025), a well-established yet orthogonal technique for compressing and accelerating neural networks, has been largely underexplored in the context of dLLMs. In the domain of auto-regressive LLMs, post-training quantization(PTQ) (Chee et al., 2024; Ashkboos et al., 2023; Tseng et al., 2024; Zhao et al., 2023) has been widely adopted to reduce the memory footprint of weights and activations, and to enable faster inference through kernel-level optimization. Yet, how well existing PTQ techniques generalize to diffusion LLMs remains an open and intriguing question.
@@ 2.1 Diffusion Language Model
Diffusion models have achieved remarkable success in image, video, and audio generation by learning to reverse a forward noise process (Jiang et al., 2025; Zhao et al., 2024a). However, applying diffusion to language generation presents unique challenges due to the discrete nature of textual data. To address this, DiffusionBERT (He et al., 2022) leverages a BERT (Devlin et al., 2019) architecture to model the reverse dynamics of a discrete diffusion process with an absorbing state, as proposed in D3PM (Austin et al., 2021a).
@@ 2.2 Network Quantization
Compared to pruning and distillation (Lin et al., 2024a; Zhang et al., 2024; Xing et al., 2025; Ye et al., 2025c, b), quantization has been extensively studied as an effective technique to compress neural networks by using low-bit representations for high-precision tensors (Zhao et al., 2024b; Xu et al., 2024; Yang et al., 2025b; Huang and Wu, 2025). Existing methods are typically categorized into two groups: post-training quantization (PTQ) and quantization-aware training (QAT). PTQ (Wu et al., 2024; Yang et al., 2024) applies quantization after model training, while QAT (Tao et al., 2022; Chen et al., 2024, 2025) incorporates quantization effects during training. Due to the high computational cost of training large language models (LLMs), PTQ has become increasingly popular for its efficiency and ability to preserve model performance without retraining (Liu et al., 2024; Dong et al., 2024). In this work, we follow this paradigm and focus on applying PTQ to dLLMs.
@@ 4.1 Experimental Setup
Model Setting Method WinoGrande PIQA ARC-C ARC-E MMLU Avg Drop

### v2

@@ Abstract
Recent advances in diffusion large language models (dLLMs) have introduced a promising alternative to autoregressive (AR) LLMs for natural language generation tasks, leveraging full attention and denoising-based decoding strategies. However, the deployment of these models on edge devices remains challenging due to their massive parameter scale and high resource demands. While post-training quantization (PTQ) has emerged as a widely adopted technique for compressing AR LLMs, its applicability to dLLMs remains largely unexplored. In this work, we present the first systematic study on quantizing diffusion-based language models. We begin by identifying the presence of activation outliers, characterized by abnormally large activation values that dominate the dynamic range. These outliers pose a key challenge to low-bit quantization, as they make it difficult to preserve precision for the majority of values. More importantly, we implement state-of-the-art PTQ methods and conduct a comprehensive evaluation across multiple task types and model variants. Our analysis is structured along four key dimensions: bit-width, quantization method, task category, and model type. Through this multi-perspective evaluation, we offer practical insights into the quantization behavior of dLLMs under different configurations. We hope our findings provide a foundation for future research in efficient dLLM deployment. Our code is publicly available at https://github.com/FelixMessi/QDLM.
@@ 1 Introduction
Current efforts toward optimizing dLLM inference have primarily focused on designing specialized key-value (KV) cache mechanisms  (Wu et al., 2025; Ma et al., 2025; Liu et al., 2025b; Wang et al., 2025). However, quantization  (Li et al., 2024; Liu et al., 2025a; Wei et al., 2025; Ye et al., 2025d), a well-established yet orthogonal technique for compressing and accelerating neural networks, has been largely underexplored in the context of dLLMs. In the domain of auto-regressive LLMs, post-training quantization(PTQ) (Chee et al., 2024; Ashkboos et al., 2023; Tseng et al., 2024; Zhao et al., 2023) has been widely adopted to reduce the memory footprint of weights and activations, and to enable faster inference through kernel-level optimization. Yet, how well existing PTQ techniques generalize to diffusion LLMs remains an open and intriguing question.
@@ 2.1 Diffusion Language Model
With the fast development of deep learning and pre-trained models (Zeng et al., 2023b, a, 2024a, 2024b, 2025; Yan et al., 2021b, 2022, a, 2023a, , 2023b, 2024b, 2024a, 2024c; Lin et al., 2025), Diffusion models have achieved remarkable success in image, video, and audio generation by learning to reverse a forward noise process (Jiang et al., 2025; Zhao et al., 2024a). However, applying diffusion to language generation presents unique challenges due to the discrete nature of textual data. To address this, DiffusionBERT (He et al., 2022) leverages a BERT (Devlin et al., 2019) architecture to model the reverse dynamics of a discrete diffusion process with an absorbing state, as proposed in D3PM (Austin et al., 2021a).
@@ 2.2 Network Quantization
Compared to pruning and distillation (Lin et al., 2024a; Zhang et al., 2024; Xing et al., 2025; Ye et al., 2025c, b), quantization has been extensively studied as an effective technique to compress neural networks by using low-bit representations for high-precision tensors (Zhao et al., 2024b; Xu et al., 2024; Yang et al., 2025b; Huang and Wu, 2025) Existing methods are typically categorized into two groups: post-training quantization (PTQ) and quantization-aware training (QAT). PTQ (Wu et al., 2024; Yang et al., 2024; Xu and Yang, 2025) applies quantization after model training, while QAT (Tao et al., 2022; Chen et al., 2024, 2025) incorporates quantization effects during training. Due to the high computational cost of training large language models (LLMs), PTQ has become increasingly popular for its efficiency and ability to preserve model performance without retraining (Liu et al., 2024; Dong et al., 2024). In this work, we follow this paradigm and focus on applying PTQ to dLLMs.
@@ 4.1 Experimental Setup
Model Setting Method WinoGrande PIQA ARC-C ARC-E MMLU 5-shot Avg Drop